### v1

# The ultimate issue error: mistaking parameters for hypotheses

Abstract
In a criminal investigation, an inferential error occurs when the probability that a suspect is the source of some evidence – such as a fingerprint – is taken as the probability of guilt. This is known as the ultimate issue error, and the same error occurs in statistical inference when the probability that a parameter equals some value is incorrectly taken to be the probability of a hypothesis. Almost all statistical inference in the social and biological sciences is subject to this error, and replacing every instance of “hypothesis testing” with “parameter testing” in these fields would more accurately describe the target of inference. The relationship between parameter values and quantities derived from them, such as p-values or Bayes factors, have no direct quantitative relationship with scientific hypotheses. Here, we describe the problem, its consequences, and suggest options for improving scientific inference.
Introduction
Suppose fingerprints are found at a crime scene and police have detained a suspect. The police hypothesise that if the suspect is guilty, their fingerprints should match those found at the scene. Upon testing, the forensic team concludes that there is a one in a million chance that the fingerprints originated from someone other than the suspect. What is the probability that the suspect is guilty? This is not a trick question about conditional probabilities, p-value interpretations, or the prosecutor’s fallacy. The answer is: we cannot determine the probability of guilt given the current information. If the crime happened in the suspect’s home, their fingerprints would be found everywhere. Therefore, finding their fingerprints at the crime scene does not imply guilt because $P(\text{fingerprint}|\text{not guilty})\approx 1$. If, however, the suspect’s fingerprints are found in the home of someone they do not know, the fingerprint evidence may be highly suggestive of guilt. This example highlights two key points. First, qualitative background information is critical for determining what a piece of evidence says about a hypothesis. Second, the degree to which the evidence implies guilt has little to do with the probability of a match; these are probabilities for different events. The probability of a fingerprint match informs the probability of guilt, but assuming that the probability of a match equals the probability of guilt is known as the ultimate issue error (Aitken et al. 2021). The “ultimate issue” is whether the suspect is guilty, and the error arises from substituting another probability; the 1:1,000,000 match probability.
A hypothesis is a testable statement or proposition. It is often expressed as a prediction based on some theory, model, or background information. A parameter is a quantitative component of a statistical model, which often represents some characteristic of a population. An example of a hypothesis is: “If this drug is effective, blood pressure in the drug group will be lower than blood pressure in the control group”. The parameter might be the mean difference in blood pressure between the two groups, denoted by $\delta$. If $\delta<0$, all we can conclude is that the hypothesis is now more plausible than before obtaining the data, but not by how much (Polya 1954). In other words, $P(\delta<0)\neq P(\text{Drug is effective})$. Observing that $\delta<0$ is a necessary but not sufficient condition for concluding that the drug is effective.
Unfortunately, the ultimate issue error is common in the social and biological sciences when testing hypotheses. Here, the ultimate issue is the probability that a hypothesis is true, but the probability calculated is whether a parameter in a statistical model equals a given value. We refer to this procedure as null hypothesis significance testing (NHST) when it is really null parameter significance testing (NPST). [^1] Further complicating matters, frequentist hypothesis testing does not directly test a hypothesis; instead, it tests if the data are inconsistent with the null hypothesis. Like a criminal investigation, what a parameter says about a hypothesis depends on qualitative background information.
We argue below that parameters and hypotheses are distinct entities and that parameter testing is usually quantitative, whereas hypothesis testing is mostly qualitative and subjective. Next, the problem of confusing the two will be described. Finally, we will discuss an approach to quantify the support for a hypothesis.
$P(\text{Parameter})\neq P(\text{Hypothesis})$
This section describes three reasons why the probability of a parameter usually does not equal the probability of a hypothesis. First, qualitative background information is always needed to interpret a parameter’s meaning in light of a hypothesis. Consider the following example. Suppose we are testing if Drug Z is effective for treating depression. We run a study with 200 subjects and find that the drug group has a 22% improvement compared to the control group, with $p=0.004$ (or a Bayes Factor of 16.7, or a posterior probability of 0.998 if you prefer a Bayesian analysis). Assume that a 22/improvement is clinically relevant. What is the probability that Drug Z is effective for treating depression?
A typical concluding statement would be “the data support the effectiveness of Drug Z for depression ($p=0.004$),” where the p-value is supposed to justify the preceding sentence, but it fails to do so. Consider the potential background information about this study and how it influences the probability of the drug’s effectiveness:
  • Double-blinded versus open-label/unblinded study.
  • Randomised controlled trial versus observational study.
  • Multi-site versus single-site study.
  • Hard primary end point (suicide) versus soft primary end point (self-reported depression).
  • A plausible mechanism of action based on preclinical data versus no known mechanism of action and no preclinical data.
  • All subjects completed the study versus twice as many subjects in the drug group were lost to follow-up.
  • Published as a registered report versus a non-registered report.
  • Funded and run by an independent academic group versus the company that developed the drug.
  • Heterogeneous subjects (both sexes, wide age range, fewer exclusion criteria) versus homogeneous subjects.
  • One author’s affiliation is a statistics department versus no authors are likely statisticians.
  • No media reports about the trial versus an article in the British Medical Journal about a whistleblower alleging “irregularities in the conduct of the trial”.
  • None of the authors had previous papers retracted versus the senior author had several papers retracted (he blames his postdoc).
  • You are familiar with the group and hold their previous work in high regard versus authors who are unknown to you.
  • Published in the New England Journal of Medicine versus a predatory journal.
  • Anonymised data and code provided in an online repository versus “data available upon request”.
  • This is the second study to obtain a positive result with this drug versus this is the first study.
Many of these points fall under the familiar categories of internal and external validity, risk of bias, and reproducibility. Few would be convinced of the drug’s effectiveness if this was an unblinded observational study using a soft endpoint with massive drop-out in the drug group, run by a lead investigator of dubious reputation who works for the company that produces the drug, and published a few days after initial submission in a predatory journal with no data provided. It does not matter how small the reported p-value or how large the Bayes factor is, there are too many problems with the study to consider the results credible. Similar to jurors weighing all the evidence to reach a verdict, researchers will take the numeric results and weigh them with relevant background information to determine if the results are convincing.
For both jurors and researchers, this is a subjective process, and is no different from what every researcher informally does when reading a paper (“this is solid” or “I doubt this will replicate”). Subjective judgements of quality are already performed when assessing the risk of bias in clinical studies (Higgins et al. 2011). Just as each jury may return a different verdict, individual researchers will uniquely weigh the above criteria, assess the degree to which the study meets or fails to meet each criterion, and come to their own conclusions. Note how the criteria can combine to influence the overall judgement. A blinded study with a soft endpoint may be convincing, as might a study without blinding but with a hard endpoint. However, an unblinded study with a soft endpoint may be too error prone to be convincing. To keep the legal analogy, having a conflict of interest (motive) and not being blinded (opportunity) might be problematic, whereas either alone might be acceptable.
Another way to see the difference between parameter testing and hypothesis testing is by noting that as the sample size increases, p-values approach zero and Bayes factors approach infinity, but we do not get correspondingly confident in the hypothesis. We can increase the sample size to move from $p=1\times 10^{-4}$ to $p=1\times 10^{-8}$, but few would consider the results four orders of magnitude more convincing. In fact, many would find it more convincing to observe two independent studies run by different groups that have $p=1\times 10^{-4}$ instead of one large study with $p=1\times 10^{-8}$.
The second reason why the probability of a parameter usually does not equal the probability of a hypothesis is when a study or experiment has multiple outcomes, and these are associated with multiple parameters. These parameters will rarely be equally distant from the null (on the appropriate scale), and may even conflict. How can the probability of a parameter equal the probability of a hypothesis when hypotheses and parameters are one-to-many? Arbitrarily defining a primary outcome to test the hypothesis only side-steps the problem.
Finally, parameters critically depend on the details of the experiment or study as well as the statistical model. For example, the drug might work better in patients with severe depression than patients with mild depression. The parameter will therefore differ based on the inclusion and exclusion criteria. Even worse is when the parameter depends on unknown or unmeasured population characteristics. For example, suppose the drug works better in patients with predominately affective symptoms (feelings of sadness or hopelessness) versus patients with predominately physiological symptoms (tiredness, lack of energy, sleep disturbances), but this is unknown to the researcher. The effect (parameter) will then largely reflect the proportion of these patient subtypes in the experiment.
$P(\text{Parameter})\approx P(\text{Hypothesis})$ only when the hypothesis is a precise numeric value, which is more common in physics and engineering. For example, based on his theory of relativity, Einstein predicted in 1911 that the sun would deflect star light by 0.83 seconds of arc, which he revised to 1.75 arcseconds in 1916 (Earman and Glymour 1980). However, even in this case, background information is important. Experimental observations that failed to support his prediction would only count against his theory if the measuring instruments were appropriate for the task, calibrated, working properly, and so on. Few if any experiments in the biological and social sciences make precise numeric predictions, and rarely even make a directional prediction as evidenced by the extensive use of two-tailed tests.
Why make the hypotheses/parameter distinction?
An injustice has occurred if a jury confuses the probability of a fingerprint match with the probability of guilt. The consequences of confusing the probability of a parameter with the probability of a hypothesis are usually less severe, but still important. Failure to distinguish between these probabilities can lead one to believe that a hypothesis has much more support than is justified. This misinterpretation may also spread to the wider public, including journalists and policymakers, where the consequences may be more severe.
In scientific publications, researchers must convince skeptical colleagues of the validity of their claims. The distinction between the probability of parameters and hypotheses emphasises the fact that small p-values alone do not provide convincing evidence. Researchers must also use appropriate methods, rule out competing explanations, ensure that the results are unbiased and therefore credible, ensure the construct is appropriately operationalised, and so on. Focusing on aspects that increase the probability of a hypothesis will lead to better experiments, compared to exclusively focusing on aspects that lead to a higher probability of rejecting a null parameter value, such as increased sample size. How do hypotheses and parameters relate to one another?
Parameters informing hypotheses
The scientific – or any other – community has not developed a general method for moving from a quantitative value for a parameter to a quantitative statement about a proposition, such as a scientific hypothesis. However, Pierce described an approach in 1878 (Peirce 2014), which has been modified and applied by many others (Good 1950; Jaynes 2003; Edwards 1992; Pardo and Allen 2007; Allen and Pardo 2019; Fairfield and Charman 2022). Essentially, it is a comparison of the relative likelihood or plausibility of the evidence ($E$) under two competing hypotheses or explanations ($H_{1}$ and $H_{0}$), expressed as a ratio of one hypothesis to another, giving a likelihood ratio (LR).
${LR=\dfrac{P(E|H_{1})}{P(E|H_{0})}.}$ (1)
Continuing our legal example, let $E$ represent the fingerprint evidence, $H_{1}$ represent the prosecution’s hypothesis that the defendant is guilty, and $H_{0}$ represent the defence’s position that the accused is not guilty and the fingerprint match is the result of chance. If the accused is guilty, we expect to find their fingerprints at the crime scene, so $P(E|H_{1})\approx 1$. If they are not guilty, the random match probability is $P(E|H_{0})=1/1,000,000$, which gives
$LR=\dfrac{P(E|H_{1})}{P(E|H_{0})}=\dfrac{1}{10^{-6}}=10^{6},$
interpreted as the evidence being one million times more likely under the prosecution’s position. However, this is not the likelihood that the accused is guilty, only the likelihood that they are the source of the fingerprint. To establish guilt, this evidence must be considered with all the other evidence, such as a lack of motive and a strong alibi, for example. Reasonable people would agree that having no motive and an alibi decreases the probability of guilt, but opinion will differ as to how much. This is the crux of the problem, and why criminal cases ask twelve jurors for their judgement.
The LR quantifies how strongly evidence supports one hypothesis or the other, but Equation 1 does not consider the other evidence required to interpret the match probability. Let’s augment Equation 1 to include all the evidence ($E^{*}$) relevant for assessing how the match probability influences the probability of guilt, as well as all the background information ($I$), which includes common-sense information such as motives increase the likelihood of guilt”
${LR^{*}=\dfrac{P(E^{*}|H_{1},I)}{P(E^{*}|H_{0},I)}.}$ (2)
The augmented likelihood ratio ($LR^{*}$) combines the quantitative random match probability defined in Equation 1 with qualitative information about motive, alibi, the probability of a laboratory error, the uncertainty about the reference population used to calculate the 1:1,000,000 number, and so on. At this point, judgement is required. It is the responsibility of the judge or jury in a criminal trial, or the editor, peer-reviewer, or other scientist when evaluating scientific findings.
Rather than dealing with $LR^{*}$, it is often more convenient to take the logarithm. By convention, a base-10 logarithm is used and the resulting log-likelihood ratio is multiplied by 10 to give decibel units: $10\,\text{Log}_{10}(LR^{*})$. This quantity is known as the weight of evidence (WoE) (Peirce 2014; Good 1950). Logarithms have several advantages. First, independent pieces of evidence can be added together to get an overall WoE. Second, a logarithmic decibel scale is more intuitive (once you are familiar with it). For example, WoE = 0 means that both hypotheses have equal support, and WoE > 0 favours $H_{1}$ while WoE < 0 favours $H_{0}$. The WoE is symmetric whereas the LR is not. For example, a likelihood ratio of 15 favours $H_{1}$ to the same degree that a likelihood ratio of 0.07 favours $H_{0}$, but this is not obvious ($0.07=1/15$). The corresponding values on a logarithmic scale are $11.810\,\text{log}_{10}(15)=$ 11.8 and $11.810\,\text{log}_{10}(1/15)=-$ 11.8, making this symmetric relationship clear. Furthermore, the difference between 0.91 and 0.99 seems larger than that between 0.99 and 0.999, yet both are 10 decibels apart. Finally, psychophysics has shown that perceptions of stimulus intensity are often proportional to the logarithm of intensity (Colman 2009), and by analogy, the degree to which beliefs should be updated is proportional to the logarithm of evidence.
Table 1 shows the relationship between the WoE, odds, and probability of hypothesis $H_{1}$ over $H_{0}$. A 10 unit increase in the WoE corresponds to a 10×10$\times$10 × increase in the odds. Three decibels makes $H_{1}$ about twice as likely as $H_{0}$, and 12 decibels corresponds to about a 0.95 probability in favour of $H_{1}$ over $H_{0}$.
Table 1: Relationship between the weight of evidence (WoE) for a hypothesis ($H$) provided by the evidence ($E$), odds, and probabilities.
WoE($H:E$) Odds Probability
0 1:1 0.5
3 2:1 0.67
6 4:1 0.8
10 10:1 0.91
12 19:1 0.95
20 100:1 0.99
30 1000:1 0.999
$P(E^{*}|H)$ in Equation 2 is the probability of the evidence given a hypothesis, but we are interested in the probability of a hypothesis given the evidence: $P(H|E^{*})$. Bayes’ Theorem is the standard way to convert from the first probability to the second, but it requires one additional component: the prior probability of each hypothesis $P(H,I)$. In a legal setting, this term captures the presumption of innocence. Combining all of these terms gives $\text{WoE}(H_{1}:E^{*},I)$, which is read as “the weight provided by $E^{*}$ and $I$ for hypothesis $H_{1}$
${\text{WoE}(H_{1}:E^{*},I)=10\,\text{log}_{10}\left[\dfrac{P(E^{*}|H_{1},I)}{P(E^{*}|H_{0},I)}\right]+10\,\text{log}_{10}\left[\dfrac{P(H_{1},I)}{P(H_{0},I)}\right].}$ (3)
The above procedure maps cleanly from a legal setting to a scientific one: $E$ is the evidence for an effect or association, and for simplicity we can take $p<0.05$ as evidence for an effect [^2] We interpret a small p-value as evidence that the experiment demonstrated an effect; we are merely using this as an indicator that “something happened”. Any criterion can be used such as Bayes Factors or posterior densities past some threshold.. $E^{*}$ includes the details of the study’s design, data collection, analysis, and potential biases as well as any conflicts of interest. $H_{0}$ and $H_{1}$ are the null and alternative hypotheses, respectively. $I$ once again is all our background information, such as “unblinded studies are more likely to be biased”.
How can we quantify these terms? $P(E^{*}|H_{0},I)$ is the probability of observing a significant result if the null hypothesis is true, which is the probability of a false positive. This quantity is usually set to $\alpha=0.05$ by the experimenter. However, the actual probability of a false positive can be larger than 0.05 because of biases, perverse incentives, questionable research practices (QRP), incompetence, or fraud. A judgement regarding how much $\alpha$ should be increased from the nominal 0.05 value can be made by anyone evaluating the experiment. If available, estimates of fraud rates or QRPs in a field can be informative, as will the specific details of the study design and execution. Note that the probability of a false positive result is not the calculated p-value – it is based on $\alpha$, and adjusted as needed based on the other evidence.
$P(E^{*}|H_{1},I)$ represents the statistical power of the experiment. This value might be reported if a power calculation was performed. Or, it might be possible to estimate the power from past studies if no information regarding the power is provided. For example, (Button et al. 2013) reported that the average power in neuroscience experiments was around 21%. The reported or estimated power provides a starting value, but it will also need to be updated to reflect the details of the experiment. Note that we are not referring to retrospective or observed power, which is calculated from the data and has a 1:1 mapping with the p-value (Hoenig and Heisey 2001; Levine and Ensom 2001; Senn 2002).
Power and sample size calculations are often simple approximations of the experiment to be conducted; they may use unrealistic effect sizes or within-group variances, and may exclude relevant aspects of the design. The actual power of our hypothetical antidepressant experiment may be lower than the calculated value due to the following reasons:
  • The dose was too low, so the effect would be small.
  • The drug was given for too short a duration before assessing its effectiveness.
  • The drug was given to some subjects who were unlikely to respond or benefit (e.g. those with severe depression). The overall effect will therefore be diluted.
  • The subjects are heterogeneous, making the outcome highly variable.
  • There is the potential for spill-over of treatment effects between groups, which makes them more alike. For example, if the treated group shares drugs with the control group.
  • The dose-response relationship is not linear (e.g. inverted “U” shape), and the drug was given at too high a dose.
  • A sub-optimal route of administration or formulation was used. For example, the best results are via injection, but it was given orally for logistical reasons.
  • The subjects included in the study have mild symptoms, and there is little room for the drug to demonstrate an improvement (floor effect).
  • A surrogate outcome for depression is used that has a weak relationship with the true outcome of interest.
  • The primary outcome had many missing values, reducing the expected sample size.
Although most people will agree that the above points reduce the power of an experiment, there will be disagreement as to how much. Judgement, once again, is required to determine as suitable value for the actual power of the experiment.
The final component of Equation 3 is the prior, which is the term to the right of the addition sign. The ratio represents the relative likelihood of the two hypotheses, before considering the data from the present experiment. This value could be based on results from past experiments, or a general assessment of the plausibility of each hypothesis. If this is the first experiment, or if we want to draw conclusions independent of previous experiments, then giving the two hypotheses equal prior plausibility will make the log of the ratio equal to zero, thus dropping this term from consideration. As a result, the WoE will only be influenced by the log-likelihood ratio.
Another way to think of the likelihood ratio is in the context of a diagnostic test, where $H_{1}$ = “person has a disease”, $H_{0}$ = “person does not have the disease”, and $E$ is a positive test result, giving
${\dfrac{P(E|H_{1})}{P(E|H_{0})}=\dfrac{P(\text{positive test | disease})}{P(\text{positive test | no disease})}=\dfrac{\text{sensitivity}}{1-\text{specificity}}=\dfrac{\text{sensitivity}}{\text{false positive rate}}.}$ (4)
The numerator is also known as the sensitivity (the proportion of positive tests among all the people with the disease), and the denominator is equal to one minus the specificity (the proportion of people who test negative among those who do not have the disease). The sensitivity and specificity are operating characteristics of a given diagnostic test. Similarly, an experiment may be characterized as having a certain sensitivity for detecting an effect (power) and the ability to correctly detect the absence of an effect (expressed as the false positive rate, or 1 - true negative rate).
Suppose, for example, that our antidepressant trial had a significant difference between groups (i.e. a positive test). The study was powered for 80%, but the drug was administered at too low a dose. Consequently, the probability of observing an effect is reduced, say the power is now only 60%. Furthermore, suppose that the researchers used $\alpha=0.05$ as their significance threshold. The informed consent form indicated that nausea was a likely side-effect of the drug, and 80% of patients on the drug indicated feeling nauseated at some point (compared with 5% of the control group). Hence, patients experiencing nausea may have concluded that they received the drug, and are now effectively unblinded. Given the subjective nature of the primary outcome, patients may expect to improve, and this expectation may bias the results such that the probability of a false positive is now increased from the nominal value of 0.05 to, say, 0.15. In the context of diagnostic tests, priors correspond to the base-rate or the prevalence of the disease in the relevant population. For this example, we decide to not use results from any previous studies and let the prior ratio equal one, and therefore log of the prior ratio equals zero and does not contribute to the result. Plugging these numbers into Equation 3 gives
$\text{WoE}(H_{1}:E^{*},I)=10\,\text{log}_{10}\left[\dfrac{0.6}{0.15}\right]+10\,\text{log}_{10}\left[1\right]=6.02+0\approx 6.$
Hence, despite a significant p-value, the WoE equals 6 (or 0.8 probability) indicating that $H_{1}$ only has modest support relative to $H_{0}$ when all the evidence about the experiment is included, and once again highlighting the difference between a parameter being unlikely given a hypothesis and the probability of the hypothesis – the ultimate issue.
If the trial returned a negative result ($p>0.05$), how likely is it that the drug is ineffective? We can perform a similar analysis, but now let $E$ = “negative result”. Using the formula for a diagnostic test we get
${\dfrac{P(E|H_{1})}{P(E|H_{0})}=\dfrac{P(\text{negative test | disease})}{P(\text{negative test | no disease})}=\dfrac{1-\text{sensitivity}}{\text{specificity}}=\dfrac{\text{false negative rate}}{\text{specificity}},}$ (5)
which gives
$\text{WoE}(H_{1}:E^{*},I)=10\,\text{log}_{10}\left[\dfrac{0.4}{0.85}\right]+10\,\text{log}_{10}\left[1\right]=-3.27+0\approx-3.$
The WoE for $H_{1}$ is -3, or equivalently, the WoE for $H_{0}$ is 3, which is negligible. We initially assumed that the hypotheses were equally likely. After the study we have only a probability of $H_{0}=0.67$. Hence, the results of this experiment are uninformative regardless of whether they are positive or negative. A positive result only provides a WoE of 6 for $H_{1}$ and a negative result gives a WoE of 3 for $H_{0}$. From these equations we can also see that if the probability of a false positive is greater than the power, a study can never provide evidence for $H_{1}$.
Discussion
Researchers are not expected to calculate the WoE in order to supplement the results of parameter testing. It is likely that researchers are unaware of flaws or biases in their studies, otherwise they would have designed them differently. Calculating the WoE is better done by scientific colleagues, and possibly peer reviewers and editors. Nevertheless, when designing a study, researchers may wish to consider the WoE calculation in order to design a more effective experiment.
Designing informative experiments
From Equation 4 and Equation 5 we can see where to focus to design informative experiments. For example, if a study is initially designed with 80% power and $\alpha=0.05$, is it better to increase power or use a stricter alpha? Based on Equation 4, we are better off decreasing the denominator (false positives) of the LR than increasing the numerator (sensitivity). The initial WoE (assuming no loss of power or bias) is $10\,\text{log}_{10}(0.8/0.05)\approx 12$. Increasing power to 95% only increases the WoE to $10\,\text{log}_{10}(0.95/0.05)\approx 12.8$, whereas decreasing false positives to $\alpha=0.01$ increases the WoE to $10\,\text{log}_{10}(0.95/0.05)\approx 19$.
Therefore, to maximize the WoE an experiment can provide, researchers should strive to keep the nominal false positive rate below $\alpha=0.05$. This should come as no surprise, since minimising bias, removing confounding, and ruling out alternative explanations are effective ways to improve an experiment.
However, when providing evidence for the null hypothesis, it is better invest effort to minimise false negatives (numerator of Equation 5), which can be verified with similar calculations.
$P(\text{Model})$ and $P(\text{Hypothesis})$
As an alternative to hypothesis testing, model comparison is another method of inference (Maxwell and Delaney 2004; Judd, McClelland, and Ryan 2009), which can be classified into two types. The first case is when two models are nested. This situation arises when one model restricts one or more parameters to specific values (usually the null value of zero or one), and the other model allows these parameters to vary and to be estimated from the data. A comparison is then made between the restricted and unrestricted models to determine if the additional flexibility of the unrestricted model produces a better fit to the data, after penalising it for its additional flexibility. This is simply another method of parameter testing since it assesses whether a parameter equals some null value or not, and the p-values will often be the same. In this case, all of the above comments are applicable.
The other type of comparison occurs when the models are not nested, that is, when one model cannot be reduced to the other by fixing parameter. An example is these two nonlinear models: $\theta_{2}(1-e^{-\theta_{1}x})$ and $\theta_{1}x/(\theta_{2}+x)$. They have qualitatively similar behaviour in that $y=0$ when $x=0$, and the functions rise to an upper asymptote as $x$ gets larger. The models provide two different stories about how the data were generated, much like the prosecution and defence provide different stories about how the evidence arose. In such cases model testing and hypothesis testing are more alike, especially when the models represent mechanistic or causal relationships and not just different empirical models. Nevertheless, background assumptions and context are still necessary (e.g. were the measurement methods sound, was the design appropriate, do the authors own shares of a company that benefits only if one model is better, etc.).
Conclusion
We have argued that in most biological and social science research the probability of a parameter cannot equal the probability of a hypothesis. First, qualitative background and contextual information is always needed to interpret the parameter value in light of the hypothesis. Second, multiple parameters associated with multiple outcomes that test the same hypothesis are rarely equal and may even conflict. Third, the value of a parameter (and its corresponding p-value or Bayes factor) often depends on unknown and unmeasured characteristics of the population or experiment.
Hence, statements such as “the treatment is effective ($p<0.05$)” makes little sense. Expanded, this statement means “we conclude that the treatment is effective because $p<0.05$”, but this is non sequitur and potentially misleading. We can only conclude that “the plausibility of the treatment being effective has increased because the data are unlikely, given the null value of a parameter in a statistical model.” The extent to which the plausibility has increased will depend on the study’s ability to detect true effects (power) and to avoid false positives, which can be quantified and represented as a WoE.
The proposed approach relies on subjective judgement to weigh evidence and interpret results, and interpretations may differ among researchers. However, this is only a formalization of the process most scientists use when reading a manuscript. A formalisation can help identify where the problem(s) are, and if one is willing to put numbers to actual power and the probability of a false positive, a quantitative estimate of the support for one hypothesis over another can be derived. In addition, formalisation can help resolve disagreements by breaking a complex inferential task into more manageable components that are evaluated separately, and then combined for a final WoE. Using this approach, for example, one researcher might conclude that $H_{1}$ has little support due to the authors’ conflict of interest, whereas another might conclude that the lack of blinding and randomisation undermines the study. This process is no different from peer review or an editorial decision to accept or reject a manuscript based on the quality of the research.
WoE provides a structured approach, but there is a risk that its subjectivity and flexibility could be exploited to justify biased interpretations, particularly in contentious or high-stakes research areas. For instance, senior scientists entrenched in the dominant paradigm could use the WoE approach to dismiss novel and promising research avenues. The WoE method, however, promotes a more comprehensive and transparent integration of evidence, which enables critical scrutiny of both data and underlying assumptions, and which can ultimately lead to more robust and credible conclusions.

### v2

@@
-The ultimate issue error: mistaking parameters for hypotheses
+The ultimate issue error in scientific inference: mistaking parameters for hypotheses
 Abstract
-In a criminal investigation, an inferential error occurs when the probability that a suspect is the source of some evidence – such as a fingerprint – is taken as the probability of guilt.
-This is known as the ultimate issue error, and the same error occurs in statistical inference when the probability that a parameter equals some value is incorrectly taken to be the probability of a hypothesis.
-Almost all statistical inference in the social and biological sciences is subject to this error, and replacing every instance of “hypothesis testing” with “parameter testing” in these fields would more accurately describe the target of inference.
-The relationship between parameter values and quantities derived from them, such as p-values or Bayes factors, have no direct quantitative relationship with scientific hypotheses.
-Here, we describe the problem, its consequences, and suggest options for improving scientific inference.
+Statistical inference often conflates the probability of a parameter with the probability of a hypothesis, a critical misunderstanding termed the ultimate issue error.
+This error is pervasive across the social, biological, and medical sciences, where null hypothesis significance testing (NHST) is mistakenly understood to be testing hypotheses rather than evaluating parameter estimates.
+Here, we advocate for using the Weight of Evidence (WoE) approach, which integrates quantitative data with qualitative background information for more accurate and transparent inference.
+Through a detailed example involving the relationship between vitamin D (25-hydroxy vitamin D) levels and COVID-19 risk, we demonstrate how WoE quantifies support for hypotheses while accounting for study design biases, power, and confounding factors.
+These findings emphasise the necessity of combining statistical metrics with contextual evaluation.
+This offers a structured framework to enhance reproducibility, reduce false interpretations, and foster robust scientific conclusions across disciplines.
 Introduction
@@
 Observing that $\delta<0$ is a necessary but not sufficient condition for concluding that the drug is effective.
-Unfortunately, the ultimate issue error is common in the social and biological sciences when testing hypotheses.
+Unfortunately, the ultimate issue error is common in the social, biological, and medical sciences when testing hypotheses.
 Here, the ultimate issue is the probability that a hypothesis is true, but the probability calculated is whether a parameter in a statistical model equals a given value.
@@
 [^1] Further complicating matters, frequentist hypothesis testing does not directly test a hypothesis; instead, it tests if the data are inconsistent with the null hypothesis.
-Like a criminal investigation, what a parameter says about a hypothesis depends on qualitative background information.
+The relationship between parameter values and quantities derived from them, such as p-values or Bayes factors, have no direct quantitative relationship with scientific hypotheses.
 We argue below that parameters and hypotheses are distinct entities and that parameter testing is usually quantitative, whereas hypothesis testing is mostly qualitative and subjective.
 Next, the problem of confusing the two will be described.
-Finally, we will discuss an approach to quantify the support for a hypothesis.
+Finally, an approach to quantify the support for a hypothesis will be presented.
 $P(\text{Parameter})\neq P(\text{Hypothesis})$
@@
 We run a study with 200 subjects and find that the drug group has a 22% improvement compared to the control group, with $p=0.004$ (or a Bayes Factor of 16.7, or a posterior probability of 0.998 if you prefer a Bayesian analysis).
-Assume that a 22/improvement is clinically relevant.
+Assume that a 22% improvement is clinically relevant.
 What is the probability that Drug Z is effective for treating depression?
-A typical concluding statement would be “the data support the effectiveness of Drug Z for depression ($p=0.004$),” where the p-value is supposed to justify the preceding sentence, but it fails to do so.
+A typical concluding statement in a manuscript would be “the data support the effectiveness of Drug Z for depression ($p=0.004$),” where the p-value is supposed to justify the preceding sentence, but it fails to do so.
 Consider the potential background information about this study and how it influences the probability of the drug’s effectiveness:
@@
   • Randomised controlled trial versus observational study.
-  • Multi-site versus single-site study.
   • Hard primary end point (suicide) versus soft primary end point (self-reported depression).
@@
   • Funded and run by an independent academic group versus the company that developed the drug.
-  • Heterogeneous subjects (both sexes, wide age range, fewer exclusion criteria) versus homogeneous subjects.
-  • One author’s affiliation is a statistics department versus no authors are likely statisticians.
-  • No media reports about the trial versus an article in the British Medical Journal about a whistleblower alleging “irregularities in the conduct of the trial”.
   • None of the authors had previous papers retracted versus the senior author had several papers retracted (he blames his postdoc).
-  • You are familiar with the group and hold their previous work in high regard versus authors who are unknown to you.
   • Published in the New England Journal of Medicine versus a predatory journal.
-  • Anonymised data and code provided in an online repository versus “data available upon request”.
-  • This is the second study to obtain a positive result with this drug versus this is the first study.
+  • It’s the second study to obtain a positive result with this drug versus it’s the first study.
 Many of these points fall under the familiar categories of internal and external validity, risk of bias, and reproducibility.
-Few would be convinced of the drug’s effectiveness if this was an unblinded observational study using a soft endpoint with massive drop-out in the drug group, run by a lead investigator of dubious reputation who works for the company that produces the drug, and published a few days after initial submission in a predatory journal with no data provided.
+Few would be convinced of the drug’s effectiveness if this was an unblinded observational study using a soft endpoint with massive drop-out in the drug group, run by a lead investigator of dubious reputation who works for the company that produces the drug, and published a few days after initial submission in a predatory journal.
 It does not matter how small the reported p-value or how large the Bayes factor is, there are too many problems with the study to consider the results credible.
@@
 For both jurors and researchers, this is a subjective process, and is no different from what every researcher informally does when reading a paper (“this is solid” or “I doubt this will replicate”).
-Subjective judgements of quality are already performed when assessing the risk of bias in clinical studies (Higgins et al.
+Some may think that subjective judgements have no place in science, but they are unavoidable, and they are already performed when assessing the risk of bias in clinical studies (Higgins et al.
 2011).
-Just as each jury may return a different verdict, individual researchers will uniquely weigh the above criteria, assess the degree to which the study meets or fails to meet each criterion, and come to their own conclusions.
 Note how the criteria can combine to influence the overall judgement.
 A blinded study with a soft endpoint may be convincing, as might a study without blinding but with a hard endpoint.
-However, an unblinded study with a soft endpoint may be too error prone to be convincing.
-To keep the legal analogy, having a conflict of interest (motive) and not being blinded (opportunity) might be problematic, whereas either alone might be acceptable.
-Another way to see the difference between parameter testing and hypothesis testing is by noting that as the sample size increases, p-values approach zero and Bayes factors approach infinity, but we do not get correspondingly confident in the hypothesis.
-We can increase the sample size to move from $p=1\times 10^{-4}$ to $p=1\times 10^{-8}$, but few would consider the results four orders of magnitude more convincing.
-In fact, many would find it more convincing to observe two independent studies run by different groups that have $p=1\times 10^{-4}$ instead of one large study with $p=1\times 10^{-8}$.
-The second reason why the probability of a parameter usually does not equal the probability of a hypothesis is when a study or experiment has multiple outcomes, and these are associated with multiple parameters.
-These parameters will rarely be equally distant from the null (on the appropriate scale), and may even conflict.
-How can the probability of a parameter equal the probability of a hypothesis when hypotheses and parameters are one-to-many?
+However, an unblinded study with a soft endpoint might have a risk of bias too high to be convincing.
+The second reason why the probability of a parameter usually does not equal the probability of a hypothesis is because there is often a one-to-many relationship between hypotheses and parameters; studies often have multiple outcomes that are associated with multiple parameters.
+For example, a study may assess depression with the Hamilton Depression Ration Scale, the Montgomery-Asberg Depression Rating Scale, and a simple self-report measure, all of which will be associated with a parameter.
+These parameters will rarely be equally distant from the null (on the appropriately standardised scale), and may even conflict.
 Arbitrarily defining a primary outcome to test the hypothesis only side-steps the problem.
+If there is a need to measure multiple outcomes to test a hypothesis, it is likely because the true outcome of interest is not directly observable or is ill-defined.
+This raises issues of construct validity (how do we know we are actually measuring depression?) and content validity (how do we know we’re measuring all aspects of it?), further weakening the link between a parameter based on a measured outcome and a hypothesis.
 Finally, parameters critically depend on the details of the experiment or study as well as the statistical model.
 For example, the drug might work better in patients with severe depression than patients with mild depression.
-The parameter will therefore differ based on the inclusion and exclusion criteria.
+The parameter will therefore differ based on the study’s inclusion and exclusion criteria.
 Even worse is when the parameter depends on unknown or unmeasured population characteristics.
@@
 The effect (parameter) will then largely reflect the proportion of these patient subtypes in the experiment.
-$P(\text{Parameter})\approx P(\text{Hypothesis})$ only when the hypothesis is a precise numeric value, which is more common in physics and engineering.
-For example, based on his theory of relativity, Einstein predicted in 1911 that the sun would deflect star light by 0.83 seconds of arc, which he revised to 1.75 arcseconds in 1916 (Earman and Glymour 1980).
-However, even in this case, background information is important.
-Experimental observations that failed to support his prediction would only count against his theory if the measuring instruments were appropriate for the task, calibrated, working properly, and so on.
-Few if any experiments in the biological and social sciences make precise numeric predictions, and rarely even make a directional prediction as evidenced by the extensive use of two-tailed tests.
+The value of a parameter and its standard error (SE) depend on the statistical model in which they are embedded.
+Models may vary in their likelihood function (e.g.
+normal or Poisson) the predictor variables included, whether nonlinearities or interactions are included, and so on.
+Several models may be reasonable, and the parameter and/or its SE will depend on the selected model.
+Therefore, no unique mapping between a parameter and a hypothesis exists.
+Even for simple experiments such as our drug example, whether the drug is effective could be defined as the mean difference between groups, the median difference, the ratio of means or medians, or something else.
+The probability of a parameter only approximates the probability of a hypothesis when the hypothesis is a precise numeric value, which is more common in physics and engineering.
+For example, based on his theory of relativity, Einstein predicted that the sun would deflect star light by 1.75 seconds of arc (Earman and Glymour 1980).
+This precisely hypothesised value was then compared with experimental observations and found to be in accordance with them.
+However, even in this case, background information was important.
+Experimental observations that failed to support Einstein’s prediction would only count against his theory if the measuring instruments were appropriate for the task, were calibrated and working properly on the day, and so on.
+Few if any experiments in the biological and social sciences make precise numeric predictions, and rarely even make a directional prediction, as evidenced by the extensive use of two-tailed tests.
 Why make the hypotheses/parameter distinction?
-An injustice has occurred if a jury confuses the probability of a fingerprint match with the probability of guilt.
-The consequences of confusing the probability of a parameter with the probability of a hypothesis are usually less severe, but still important.
 Failure to distinguish between these probabilities can lead one to believe that a hypothesis has much more support than is justified.
-This misinterpretation may also spread to the wider public, including journalists and policymakers, where the consequences may be more severe.
+This misinterpretation can also spread to the wider public, including journalists and policymakers, where the consequences can be severe.
 In scientific publications, researchers must convince skeptical colleagues of the validity of their claims.
@@
 Focusing on aspects that increase the probability of a hypothesis will lead to better experiments, compared to exclusively focusing on aspects that lead to a higher probability of rejecting a null parameter value, such as increased sample size.
-How do hypotheses and parameters relate to one another?
 Parameters informing hypotheses
-The scientific – or any other – community has not developed a general method for moving from a quantitative value for a parameter to a quantitative statement about a proposition, such as a scientific hypothesis.
+No method has been generally adopted for moving from a quantitative value for a parameter to a quantitative statement about a proposition, such as a scientific hypothesis.
 However, Pierce described an approach in 1878 (Peirce 2014), which has been modified and applied by many others (Good 1950; Jaynes 2003; Edwards 1992; Pardo and Allen 2007; Allen and Pardo 2019; Fairfield and Charman 2022).
@@
 The LR quantifies how strongly evidence supports one hypothesis or the other, but Equation 1 does not consider the other evidence required to interpret the match probability.
-Let’s augment Equation 1 to include all the evidence ($E^{*}$) relevant for assessing how the match probability influences the probability of guilt, as well as all the background information ($I$), which includes common-sense information such as motives increase the likelihood of guilt”
+Let’s augment Equation 1 to include all the evidence ($E^{*}$) relevant for assessing how the match probability influences the probability of guilt, as well as all the background information ($I$), which includes common-sense information such as “having a motive increases the likelihood of guilt”
 ${LR^{*}=\dfrac{P(E^{*}|H_{1},I)}{P(E^{*}|H_{0},I)}.}$ (2)
@@
 The WoE is symmetric whereas the LR is not.
-For example, a likelihood ratio of 15 favours $H_{1}$ to the same degree that a likelihood ratio of 0.07 favours $H_{0}$, but this is not obvious ($0.07=1/15$).
+For example, a LR of 15 favours $H_{1}$ to the same degree that a LR of 0.07 favours $H_{0}$, but this is not obvious ($0.07=1/15$).
 The corresponding values on a logarithmic scale are $11.810\,\text{log}_{10}(15)=$ 11.8 and $11.810\,\text{log}_{10}(1/15)=-$ 11.8, making this symmetric relationship clear.
@@
 Finally, psychophysics has shown that perceptions of stimulus intensity are often proportional to the logarithm of intensity (Colman 2009), and by analogy, the degree to which beliefs should be updated is proportional to the logarithm of evidence.
+There is also recent evidence that people can more easily combine information using log ratios, compared with ratios or probabilities (Juslin et al.
+2011).
 Table 1 shows the relationship between the WoE, odds, and probability of hypothesis $H_{1}$ over $H_{0}$.
@@
 Note that the probability of a false positive result is not the calculated p-value – it is based on $\alpha$, and adjusted as needed based on the other evidence.
-$P(E^{*}|H_{1},I)$ represents the statistical power of the experiment.
+$P(E^{*}|H_{1},I)$ is the probability of observing a significant result when the alternative hypothesis is true, and therefore represents the statistical power of the experiment.
 This value might be reported if a power calculation was performed.
 Or, it might be possible to estimate the power from past studies if no information regarding the power is provided.
-For example, (Button et al.
-2013) reported that the average power in neuroscience experiments was around 21%.
+For example, Button et al.
+(2013) reported that the average power in neuroscience experiments was around 21%.
 The reported or estimated power provides a starting value, but it will also need to be updated to reflect the details of the experiment.
@@
 Power and sample size calculations are often simple approximations of the experiment to be conducted; they may use unrealistic effect sizes or within-group variances, and may exclude relevant aspects of the design.
-The actual power of our hypothetical antidepressant experiment may be lower than the calculated value due to the following reasons:
+The actual power of the hypothetical antidepressant experiment may be lower than the calculated value due to the following reasons:
   • The dose was too low, so the effect would be small.
@@
   • The drug was given to some subjects who were unlikely to respond or benefit (e.g. those with severe depression). The overall effect will therefore be diluted.
-  • The subjects are heterogeneous, making the outcome highly variable.
-  • There is the potential for spill-over of treatment effects between groups, which makes them more alike. For example, if the treated group shares drugs with the control group.
   • The dose-response relationship is not linear (e.g. inverted “U” shape), and the drug was given at too high a dose.
@@
 Although most people will agree that the above points reduce the power of an experiment, there will be disagreement as to how much.
-Judgement, once again, is required to determine as suitable value for the actual power of the experiment.
+Judgement, once again, is required to determine a suitable value for the actual power of the experiment.
 The final component of Equation 3 is the prior, which is the term to the right of the addition sign.
@@
 As a result, the WoE will only be influenced by the log-likelihood ratio.
-Another way to think of the likelihood ratio is in the context of a diagnostic test, where $H_{1}$ = “person has a disease”, $H_{0}$ = “person does not have the disease”, and $E$ is a positive test result, giving
+Another way to think of the LR is in the context of a diagnostic test, where $H_{1}$ = “person has the disease”, $H_{0}$ = “person does not have the disease”, and $E$ is a positive test result, giving
 ${\dfrac{P(E|H_{1})}{P(E|H_{0})}=\dfrac{P(\text{positive test | disease})}{P(\text{positive test | no disease})}=\dfrac{\text{sensitivity}}{1-\text{specificity}}=\dfrac{\text{sensitivity}}{\text{false positive rate}}.}$ (4)
@@
 Similarly, an experiment may be characterized as having a certain sensitivity for detecting an effect (power) and the ability to correctly detect the absence of an effect (expressed as the false positive rate, or 1 - true negative rate).
-Suppose, for example, that our antidepressant trial had a significant difference between groups (i.e.
+Suppose, for example, that the antidepressant trial had a significant difference between groups (i.e.
 a positive test).
@@
 In the context of diagnostic tests, priors correspond to the base-rate or the prevalence of the disease in the relevant population.
-For this example, we decide to not use results from any previous studies and let the prior ratio equal one, and therefore log of the prior ratio equals zero and does not contribute to the result.
+For this example, results from previous studies are not used.
+The prior ratio therefore equals one, and the log of the prior ratio equals zero and does not contribute to the result.
 Plugging these numbers into Equation 3 gives
 $\text{WoE}(H_{1}:E^{*},I)=10\,\text{log}_{10}\left[\dfrac{0.6}{0.15}\right]+10\,\text{log}_{10}\left[1\right]=6.02+0\approx 6.$
-Hence, despite a significant p-value, the WoE equals 6 (or 0.8 probability) indicating that $H_{1}$ only has modest support relative to $H_{0}$ when all the evidence about the experiment is included, and once again highlighting the difference between a parameter being unlikely given a hypothesis and the probability of the hypothesis – the ultimate issue.
+Hence, despite a significant p-value, the WoE equals 6 (or 0.8 probability) indicating that $H_{1}$ only has modest support relative to $H_{0}$ when all the evidence about the experiment is included.
+This once again highlights the difference between a parameter being unlikely given a hypothesis and the probability of the hypothesis – the ultimate issue.
 If the trial returned a negative result ($p>0.05$), how likely is it that the drug is ineffective?
@@
 From these equations we can also see that if the probability of a false positive is greater than the power, a study can never provide evidence for $H_{1}$.
+Real world example: COVID-19 and vitamin D
+Using United Kingdom (UK) Biobank data, Hastie et al.
+(2020) examined the relationship between blood vitamin D levels (25-hydroxy vitamin D), and the risk of COVID-19 infection.
+Complete data were available for 348,598 participants.
+In this data set, 2,724 COVID-19 tests were conducted on 1,474 individuals, and 449 had a positive COVID-19 test.
+This group was compared to the 1,025 people who tested negative, and the 347,124 people who were not tested.
+The outcome was confirmed COVID-19 infection, defined as at least one positive test result.
+The main finding was that Vitamin D was associated with COVID-19 infection in a univariate logistic regression model (odds ratio = 0.99; 95% CI 0.99 to 0.999; p = 0.013), but not after adjustment for confounders (OR = 1.00; 95% CI = 0.998 to 1.01; p = 0.208).
+The author’s conclusion was “Our findings do not support a potential link between vitamin D concentrations and risk of COVID-19 infection…”.
+This paper has been cited nearly 600 times and referred to as “credible evidence” by factcheck.org that low vitamin D is not associated with a higher risk of COVID-19 infection.
+[^3] https://www.factcheck.org/2020/06/does-vitamin-d-protect-against-covid-19/
+Using the above approach (e.g.
+Equation 5), we can assess whether this study provides strong evidence for a lack of an association between vitamin D levels and COVID-19 infection and whether the conclusion is justified.
+Since we are only interested in quantifying the evidence in this study, prior information will not be included.
+This means the two hypotheses will be given equal initial probabilities of 0.5.
+The basic equation is then
+$\dfrac{P(E|H_{1})}{P(E|H_{0})}=\dfrac{P(\text{negative result | association exists })}{P(\text{negative result | no association})}.$
+We can initially take the denominator to be $0.951-\alpha=1-0.05=$ 0.95 (although this was not explicitly stated by the researchers).
+However, the probability of a false positive is slightly higher because the COVID-19 positive and negative groups differed on almost all baseline variables (Table 1 in Hastie et al.
+(2020)).
+Although these were adjusted for in the analysis, there may be residual confounding due to nonlinear relationships or interactions between variables.
+In addition, several analyses were conducted using univariate or multivariate models and treating vitamin D levels as continuous or binary.
+Let us modestly increase the probability of a false positive from 0.05 to 0.1, and hence the denominator will be $1-0.1=0.9$
+The numerator is of greater interest because the study was negative, and we want to know the probability of a negative result if in fact there is an association.
+An initial value for the numerator is 1 - power.
+Hastie et al.
+(2020) did not provide a power calculation, but we can get an approximate estimate of the power based on the sample size (N=348,598), the total number of positive cases (449), and the statistical model (logistic regression).
+For the power calculation, assume that vitamin D level is the only predictor and that it is a binary high/low value.
+The outcome is testing positive or negative for COVID-19.
+We need to choose a clinically meaningful effect size, and we can use an odds ratio of 0.80, which corresponds to a difference of 50 cases between the high and low vitamin D groups.
+The power for this study is only 65% [^4] R code for the calculation is provided in the supplementary material., and therefore the numerator is $1-0.65=0.35$.
+Entering these numbers into our equation gives a value of
+$\text{WoE}(H_{1}:E^{*},I)=10\,\text{log}_{10}\left[\dfrac{0.35}{0.90}\right]+10\,\text{log}_{10}\left[1\right]=-4.1.$
+In other words, if we started with both hypotheses having an equal probability of 0.5, after this study, the probability of no association ($H_{0}$) has increased to only 0.72, which is a modest change and hardly convincing.
+The situation is much worse however.
+This calculation is based only on the study design and does not account for two critical issues that considerably reduce the power.
+First, vitamin D levels were not measured just before participants were exposed to the virus, but 10-14 years earlier!
+The measured vitamin D levels are therefore a noisy proxy for the actual vitamin D levels.
+Jorde et al.
+(2010) estimated that vitamin D measurements taken 14 years apart have a Pearson correlation of 0.42 to 0.52 (depending on the method used to adjust for seasonal variation), and Meng et al.
+(2012) estimated a Spearman correlation of 0.61 when measurements were 5.1 years apart.
+The additional noise introduced by using a proxy variable will obscure any true association between vitamin D levels and COVID-19 infection.
+The second major issue is that the COVID-19 negative group were comprised of 1,025 people who tested negative, and 347,124 people who were not tested but just assumed to be negative.
+Many of these people likely had COVID-19 but were not tested if they were not hospitalised, and Davies, Mazess, and Benskin (2021) estimate that there could have been nine times as many positive people in the “negative” group than in the “positive” group.
+This introduces considerable misclassification error, which will bias results towards the null hypotheses of no association.
+Taken together, these issues greatly reduce power from the already low value of 65%.
+Simulations could be conducted if an accurate power estimate is important, but let’s estimate the power to be 20%, which gives $1-0.2=0.8$ for the numerator of the above formula, and therefore $\text{WoE}(H_{1}:E^{*},I)=-0.75$.
+In other words, the probability of “no association” has moved from an initial value of 0.5 to a final value of 0.54!
+This study had additional issues that would reduce power further (Davies, Mazess, and Benskin 2021), but let’s stop here.
+In summary, this study has provided zero evidence for a lack of an association between vitamin D and COVID-19 infection.
 Discussion
@@
 This should come as no surprise, since minimising bias, removing confounding, and ruling out alternative explanations are effective ways to improve an experiment.
-However, when providing evidence for the null hypothesis, it is better invest effort to minimise false negatives (numerator of Equation 5), which can be verified with similar calculations.
-$P(\text{Model})$ and $P(\text{Hypothesis})$
-As an alternative to hypothesis testing, model comparison is another method of inference (Maxwell and Delaney 2004; Judd, McClelland, and Ryan 2009), which can be classified into two types.
-The first case is when two models are nested.
-This situation arises when one model restricts one or more parameters to specific values (usually the null value of zero or one), and the other model allows these parameters to vary and to be estimated from the data.
-A comparison is then made between the restricted and unrestricted models to determine if the additional flexibility of the unrestricted model produces a better fit to the data, after penalising it for its additional flexibility.
-This is simply another method of parameter testing since it assesses whether a parameter equals some null value or not, and the p-values will often be the same.
-In this case, all of the above comments are applicable.
-The other type of comparison occurs when the models are not nested, that is, when one model cannot be reduced to the other by fixing parameter.
-An example is these two nonlinear models: $\theta_{2}(1-e^{-\theta_{1}x})$ and $\theta_{1}x/(\theta_{2}+x)$.
-They have qualitatively similar behaviour in that $y=0$ when $x=0$, and the functions rise to an upper asymptote as $x$ gets larger.
-The models provide two different stories about how the data were generated, much like the prosecution and defence provide different stories about how the evidence arose.
-In such cases model testing and hypothesis testing are more alike, especially when the models represent mechanistic or causal relationships and not just different empirical models.
-Nevertheless, background assumptions and context are still necessary (e.g.
-were the measurement methods sound, was the design appropriate, do the authors own shares of a company that benefits only if one model is better, etc.).
+However, when providing evidence for the null hypothesis, it is better to minimise false negatives (numerator of Equation 5), which can be verified with similar calculations.
+Evaluating manuscripts and published evidence
+The proposed approach relies on subjective judgement to weigh evidence and interpret results, and interpretations may differ among researchers.
+However, this is only a formalization of the process most scientists use when reading a manuscript.
+Formalisation can help identify where the problem(s) are, and if one is willing to put numbers to actual power and the probability of a false positive, a quantitative estimate of the support for one hypothesis over another can be derived.
+This process is no different from what peer reviewers and editors informally do when deciding to accept or reject a manuscript based on the quality of the research.
+It is an empirical question whether such an approach would improve peer-review and the editorial assessment of manuscripts, perhaps with the aid of a checklist to standardise the criteria considered.
+Similarly, published papers can be evaluated with a WoE approach, as was done for the vitamin D and COVID-19 example.
+Papers such as Hastie et al.
+(2020) provide nearly zero information for a vitamin D/COVID-19 association and can therefore be ignored, and certainly not included in meta-analyses, where they can bias the results.
 Conclusion
@@
 Second, multiple parameters associated with multiple outcomes that test the same hypothesis are rarely equal and may even conflict.
-Third, the value of a parameter (and its corresponding p-value or Bayes factor) often depends on unknown and unmeasured characteristics of the population or experiment.
+Third, the value of a parameter (and its corresponding p-value or Bayes factor) often depends on unknown and unmeasured characteristics of the population or experiment as well as the statistical model.
 Hence, statements such as “the treatment is effective ($p<0.05$)” makes little sense.
@@
 We can only conclude that “the plausibility of the treatment being effective has increased because the data are unlikely, given the null value of a parameter in a statistical model.” The extent to which the plausibility has increased will depend on the study’s ability to detect true effects (power) and to avoid false positives, which can be quantified and represented as a WoE.
-The proposed approach relies on subjective judgement to weigh evidence and interpret results, and interpretations may differ among researchers.
-However, this is only a formalization of the process most scientists use when reading a manuscript.
-A formalisation can help identify where the problem(s) are, and if one is willing to put numbers to actual power and the probability of a false positive, a quantitative estimate of the support for one hypothesis over another can be derived.
-In addition, formalisation can help resolve disagreements by breaking a complex inferential task into more manageable components that are evaluated separately, and then combined for a final WoE.
-Using this approach, for example, one researcher might conclude that $H_{1}$ has little support due to the authors’ conflict of interest, whereas another might conclude that the lack of blinding and randomisation undermines the study.
-This process is no different from peer review or an editorial decision to accept or reject a manuscript based on the quality of the research.
 WoE provides a structured approach, but there is a risk that its subjectivity and flexibility could be exploited to justify biased interpretations, particularly in contentious or high-stakes research areas.
 For instance, senior scientists entrenched in the dominant paradigm could use the WoE approach to dismiss novel and promising research avenues.
-The WoE method, however, promotes a more comprehensive and transparent integration of evidence, which enables critical scrutiny of both data and underlying assumptions, and which can ultimately lead to more robust and credible conclusions.
+The WoE method, however, promotes a more comprehensive and transparent evaluation of evidence, which enables critical scrutiny of both data and underlying assumptions, ultimately leading to more robust and credible conclusions.